You are tasked with helping a cyclist improve their linear speed. Doing a thorough analysis, you determine these current specs.:

Wheel Size: 650 millimeters

Crank Speed: 754 radians/min

Gear ratio: 1.25, meaning wheel rotates @ 1.25 x Crank Speed

Linear Speed: 10.21 meters/sec

The simplest modification you come up with is to increase the person’s Gear Ratio from 1.25 to 2.0 and decrease their cadence from 120 to 80 rpm. What would the cyclist’s new Linear Speed be in m/sec?
Answer #1

Wheel Size: 650 millimeters= 0.65 m

decrease their cadence from 120 to 80 rpm ..

so new crank speed = 80 rpm = 80* 2π= 502.4 radians/min

increase the person’s Gear Ratio from 1.25 to 2.0..

Gear ratio: 2.0

New Linear Speed = Gear ratio* Wheel Size *crank speed / 60

= (2.0)*(0.65)*(502.4) / 60 =10.885 m/s